OpenStudy (anonymous):

Line M is represented by the following equation: x − y = 8 Which equation completes the system that is satisfied by the solution (18, 10)? 2x − y = 26 x + y = 18 2x − 2y = 36 x − y = −28

OpenStudy (anonymous):

@angel_mejia

Vocaloid (vocaloid):

(18,10) tells us that x = 18 and y = 10 so, go through each answer choice one by one, and plug in x = 18 and y = 10, and see which one is true

OpenStudy (anonymous):

c? @Vocaloid

Vocaloid (vocaloid):

nope, check your math again

OpenStudy (anonymous):

b?

Vocaloid (vocaloid):

please do not guess, go through each answer choice one by one and plug in the numbers

Vocaloid (vocaloid):

start with the first choice: 2x − y = 26, true or false?

OpenStudy (anonymous):

I'm having really hard time and have a huge time crunch

Vocaloid (vocaloid):

again, start with the first choice... 2x - y = ?

OpenStudy (anonymous):

i don't even know where to start :c

Vocaloid (vocaloid):

x = 18 y = 10 2x = ?

OpenStudy (anonymous):

36?

Vocaloid (vocaloid):

right, so 2x - y = ?

OpenStudy (anonymous):

26

Vocaloid (vocaloid):

right. the first choice says that 2x − y = 26, so this is our answer

OpenStudy (anonymous):

thank u sm,can u do another plz?

Vocaloid (vocaloid):

sure

OpenStudy (anonymous):

give me one sec

OpenStudy (anonymous):

Line M is represented by the following equation: x + y = −1 What is most likely the equation for line P so the set of equations has infinitely many solutions? 2x + 2y = 2 2x + 2y = 4 2x + 2y = −2 x − y = 1

Vocaloid (vocaloid):

hint: multiply x + y = -1 by 2 on both sides

OpenStudy (anonymous):

im so confused ;c

Vocaloid (vocaloid):

just multiply everything by 2.... x + y = -1

OpenStudy (anonymous):

so x+y=-1 multiplied by 2 would be 2x+2y=-2?

Vocaloid (vocaloid):

yes, and that's your answer

OpenStudy (anonymous):

thank u! one more?

Vocaloid (vocaloid):

sure..

OpenStudy (anonymous):

Two lines, A and B, are represented by the following equations: Line A: y = x − 1 Line B: y = −3x + 11 Which of the following options shows the solution to the system of equations and explains why? (3, 2), because the point does not lie on any axis (3, 2), because one of the lines passes through this point (3, 2), because the point lies between the two axes (3, 2), because both lines pass through this point

Vocaloid (vocaloid):

hint: first multiply line A by 3

OpenStudy (anonymous):

y=x -3?

OpenStudy (anonymous):

@Vocaloid

Vocaloid (vocaloid):

not quite, you need to multiply EVERYTHING by 3

OpenStudy (anonymous):

3y=3x-3

Vocaloid (vocaloid):

right, next step: line A: 3y = 3x - 3 line B: y = -3x + 11 now add line A + line B

OpenStudy (anonymous):

3y=0x+14

Vocaloid (vocaloid):

nope, check your math again... 3y + y = ?

OpenStudy (anonymous):

im not sure

Vocaloid (vocaloid):

3y + y = __y fill in the blank

OpenStudy (anonymous):

4y

OpenStudy (anonymous):

what does y equal?

Vocaloid (vocaloid):

we're about to find out! now, when we add line A and line B together, we get 4y on the left side. and -3+11 on the right side. so 4y = -3+11

Vocaloid (vocaloid):

-3+11 = ?

OpenStudy (anonymous):

8

Vocaloid (vocaloid):

right. so 4y = 8 what is y?

OpenStudy (anonymous):

4?

Vocaloid (vocaloid):

4y = 8 divide both sides by 4, what do you get?

OpenStudy (anonymous):

1,2

Vocaloid (vocaloid):

right, so we get y = 2

Vocaloid (vocaloid):

now we go to our second equation: y = -3x + 11 now we plug in y = 2 2 = -3x + 11 what is x?

OpenStudy (anonymous):

I'm sorry i have to go to lunch with my fan and this is my last question, can i plz just have the answer